OpenStudy (anonymous):

(8th grade Algebra) Medal & Fan :) .. please help asap..

OpenStudy (anonymous):

What is the solution to the equation 4x + 2(x – 3) = 3x + x – 12? –3 –1 1 3

OpenStudy (aveline):

Let's simplify the equation using the distributive property: 4x+2x-6=3x+x-12 Next, combine x's: 6x-6=4x-12 You should be able to solve from here.

OpenStudy (anonymous):

THANK YOU! I was so confused with the x on the other side of the = sign! Thanks so much, what a big help!!!

OpenStudy (anonymous):

you now have 14 fans :3 congrats!